OpenStudy (anonymous):

A lead block has a mass of 250 grams. The block is initially at a room temperature of 22 Celsius. The melting temperature of lead is 328 Celsius. How much energy is required to warm the lead block to its melting temperature? It's grade 11 physics.

OpenStudy (theeric):

Hi! For this, you want to know the "specific heat" of lead. That is, how much energy do you need so that you cause a given temperature change for a given amount of mass. http://quiz2.chem.arizona.edu/preproom/Demo%20Files/specific_heat_of_lead.htm This link says that the specific heat of lead is \(0.16\ \rm\dfrac J {g\ C^\circ} \) You just make the units correct, now. Multiply by the grams and the Celsius degrees to be left with the Joules! So \(E={\rm specific\ heat}\times m\times \Delta T\)